Write the equation of the line passing through the given pair of points. Write your result in slope-intercept form.
(-1,3) and (4,-2)

Answer by checkley71(8403)

Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
SLOPE=(-2-3)/(4+1)=-5/5=-1
NOW REPLACE THE X & Y TERMS WITH ONE SET OF POINTS (-1,3)
3=-1*-1+b
3=1+b
b=3-1
b=2
Y=-X+2 ANSWER.